OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0). y − 4 = −one half(x + 6) y − 4 = 2(x + 6) y + 6 = −one half(x − 4) y + 6 = 2(x − 4)

OpenStudy (anonymous):

First you need to plug in each equation the X point you get the Y point solve it until one works Let's take the first equation works for the first point, try the second in it now 0-4=-1/2(2+6) -4=-1/2(8) -4=-4

alones (alones):

I'm pretty sure it's y − 4 = −1/2(x + 6) since loking at his equation

OpenStudy (anonymous):

Okay thank you